OpenStudy (anonymous):

If x* 2-9+11=23 what is x

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

I will give medal and fan

OpenStudy (misty1212):

what does x*2 mean?

OpenStudy (anonymous):

X times 2

OpenStudy (misty1212):

hmm really? \[x\times 2-9+11=23\] like that? seems unlikely but i guess it could be true

OpenStudy (anonymous):

What is x

OpenStudy (misty1212):

most people write \[2x+2=23\]

OpenStudy (misty1212):

can you post a screen shot? i am still not sure it is what you wrote if it is a) subtract 2 then b) divide by 2